# Adiabatic Bottlenecks in Quantum Annealing and Nonequilibrium Dynamics of Paramagnons

### v1

@@ Abstract
The correspondence between long-range interacting quantum spin glasses and combinatorial optimization problems underpins the physical motivation for adiabatic quantum computing. On one hand, in disordered (quantum) spin systems, the focus is on exact methods such as the replica trick that allow the calculation of system quantities in the limit of infinite system and ensemble size. On the other hand, when solving a given instance of an optimization problem, disorder-averaged quantities are of no relevance, as one is solely interested in instance-specific, finite-size properties, in particular the true solution. Here, we apply the nonequilibrium Green-function formalism to the spin coherent-state path integral to obtain the statistical fluctuations and the collective-excitation spectrum along the annealing path. For the example of the quantum Sherrington-Kirkpatrick spin glass, by comparing to extensive numerically exact results, we show that this method provides access to the instance-specific bottlenecks of the annealing protocol.
@@ I Introduction
This leads us back to the “most fundamental problem” of quantum annealing [4]: how to obtain some information about the size and, in particular, the onset of the minigap without solving the problem. Here, we demonstrate that the semi-classical fluctuations around the spin mean-field introduced previously [10] can serve as an indicator of the critical point. Physically, these fluctuations describe the dynamics of paramagnons away from equilibrium. They are obtained by solving the equations of motion (II.4) of the nonequilibrium Green functions [11, 12], which in turn are derived from the Gaussian approximation of the spin coherent-state path integral [13, 11, 10].
@@
To obtain numerically exact data as a comparison for our semi-classical methods, we perform very extensive simulations. Specifically, we diagonalize a large number of random SK instances at different system sizes and retain only the subset with the smallest minimal spectral gaps. The instances in this subset then all contain bottlenecks of varying criticality when annealing the transverse field. The details of these data are provided in Appendix B. Interestingly, even within this critical subset, there remain instances for which the adiabatic mean-field evolution actually finds the true ground state. Our main analysis, however, is then further restricted to those critical instances for which the mean-field evolution fails. This appears natural since we want to investigate problems hard for quantum annealing – but any instance solved via mean-field can be considered easy. Importantly, we find that the information provided by our semi-classical approximation about the critical point of the annealing schedule is robust: the method works even when the mean-field algorithm fails, such that we succeed in showing that the alignment between the critical fluctuations and the adiabatic bottleneck is preserved also on average (Fig. 8).
@@ II Methods
The couplings $J_{ij}$ and the local magnetic fields $h_{i}$ are i.i.d. standard normal random variables (s. Appendix B). We set $\Gamma=1$ everywhere, i.e. we consider frequencies to be given in units of $\Gamma$. The adiabatic time is confined to $0\leq t\leq T_{f}$, which makes it natural to also introduce the scaled time $s(t)=t/T_{f}$, $s(t)\in[0,1]$. This results in the adiabatic Hamiltonian
@@ II.1 Mean-Field Dynamics
By inspecting many random instances of $h_{i}$ and $J_{ij}$, one finds that the trajectories typically obey
@@ II.2 Maximally Frustrated Spins
As described vividly by Mézard, Parisi, and Virasoro in the introduction to Ref. [1], frustration is an important and ubiquitous phenomenon. Also in the present context, the tug of opposing forces can have dramatic consequences, leading to strongly frustrated spins that cannot “decide” whether to point to the north or south pole of their Bloch spheres. We illustrate this in Fig. 4, where one spin fluctuates between the hemispheres around the equator. For our random data set detailed in Appendix B, this can be observed to be quite characteristic of hard instances going through a critical point or the finite-size analog of a first-order phase transition, the onset of which finally forces the frustrated spin either north or south.
@@ II.3 Gaussian Fluctuations
The Gaussian fluctuations are bosonic and thus implicitly define creation and annihilation operators $\hat{\eta}_{i}^{{\dagger}}$, $\hat{\eta}_{i}$ for quasi-particles one may imagine as paramagnons. In terms of these operators, we can introduce the so-called greater and lesser Green functions [23, 24]
$G_{ij}^{>}(t,t^{\prime})$ $=-\mathrm{i}\langle\hat{\eta}_{i}(t)\hat{\eta}_{j}^{{\dagger}}(t^{\prime})\rangle,$ (23a)
$G_{ij}^{<}(t,t^{\prime})$ $=-\mathrm{i}\langle\hat{\eta}_{j}^{{\dagger}}(t^{\prime})\hat{\eta}_{i}(t)\rangle.$ (23b)
The diagonals of these two matrices have an intuitive meaning at equal times, i.e. for $t^{\prime}\to t$ (s. also Fig. 5). By letting $\mathrm{i}G_{ii}^{>}(t,t)=\mathcal{N}_{i}+1$ and $\mathrm{i}G_{ii}^{<}(t,t)=\mathcal{N}_{i}$, we can interpret $\mathcal{N}_{i}$ as the number of paramagnons (for a bosonic light field, this would just be the number of photons). The only difference between the on-site greater and lesser Green functions at equal times is then given by the zero-point contribution to $G_{ii}^{>}(t,t)$. By inspecting Eq. (20), we see that the corresponding anomalous Green functions proportional to $\langle\hat{\eta}_{i}(t)\hat{\eta}_{j}(t^{\prime})\rangle$ will also appear in the description of the problem. For completeness, note that the time-dependent Bogoliubov Hamiltonian corresponding to Eq. (20) can be written as
$\begin{split}\hat{H}(t)=\frac{1}{2}\sum_{i,j=1}^{N}\Big{[}&A_{ij}(t)\hat{\eta}_{i}^{\dagger}\hat{\eta}_{j}+A_{ji}(t)\hat{\eta}_{j}^{\dagger}\hat{\eta}_{i}\\[-5.69054pt] +&B_{ij}(t)\hat{\eta}_{i}^{\dagger}\hat{\eta}_{j}^{\dagger}+\bar{B}_{ji}(t)\hat{\eta}_{i}\hat{\eta}_{j}\Big{]}.\end{split}$ (24)
@@ II.4 Schwinger-Keldysh Formalism
$\bm{G}^{>}(t,t^{\prime})$ $=-\mathrm{i}\left\langle\begin{pmatrix}\eta^{-}(t)\\ \bar{\eta}^{-}(t)\end{pmatrix}\begin{pmatrix}\bar{\eta}^{+}(t^{\prime})&\eta^{+}(t^{\prime})\end{pmatrix}\right\rangle,$ (26d)
$\bm{G}^{<}(t,t^{\prime})$ $=-\mathrm{i}\left\langle\begin{pmatrix}\eta^{+}(t)\\ \bar{\eta}^{+}(t)\end{pmatrix}\begin{pmatrix}\bar{\eta}^{-}(t^{\prime})&\eta^{-}(t^{\prime})\end{pmatrix}\right\rangle.$ (26h)
@@ II.5 Statistical and Spectral Functions
Instead of working with the Green functions defined by Eqs. (II.3) or (II.4), respectively, it is conventional to transform to the following equivalent linear combinations,
@@
We will refer to $\bm{F}(t,t^{\prime})$ as the statistical function, while $\bm{\rho}(t,t^{\prime})$ is the so-called spectral function. Similarly to our above discussion of the equal-time functions, it now follows that $\mathrm{i}F_{ii}(t,t)=2\mathcal{N}_{i}+1$ is directly related to the paramagnon number. The spectral function, in turn, has the property $\bm{\rho}(t,t)=-\mathrm{i}\mathds{1}$, the origin of which lies in the bosonic commutation relation. An illustration of the behavior of this function away from the equal-time diagonal is provided in Fig. 5. The equation of motion of the statistical function can be derived easily by combining Eqs. (II.4) to yield
@@
Furthermore, the solution of the full nonequilibrium dynamics, as developed here, requires two further equations of motion, one for the statistical and one for the spectral function. The former can be derived from Eq. (27a) by adding the respective equations for $\bm{G}^{>}$ and $\bm{G}^{<}$; it is not relevant here since we only require the equal-time dynamics $\bm{F}(T,0)=\bm{F}(t,t)$, which, in the Gaussian case, decouples from the dynamics away from the forward time diagonal $t=t$ (s. Fig. 5). The remaining equation of motion for the spectral function can also be derived straightforwardly from Eqs. (II.4) and reads
@@ III.1 Illustration: Details of a Hard Instance
For our current example, the total EA order parameter from Eq. (12), plotted in blue in the second panel, does not show a very distinct increase around the critical point, which is due to the fact that the already localized spins have the largest contributions. This underlines the usefulness of focusing on the critical spin. Finally, consider the effective magnetization shown in the last panel of Fig. 3. Intriguingly, we perceive that our critical spin is completely frustrated during the entire evolution, an observation that we have found repeated throughout the full data set described in Appendix B. For this reason, we refer to the critical spin as maximally frustrated.
@@
to which we will refer as the localization susceptibility. It provides a joint measure of how localized and susceptible to fluctuations the maximally frustrated spin is at any given time. The resulting curve for our current example is presented in Fig. 7 alongside the adiabatic ratio $g_{1}(s)$ as defined in Eq. (3). The peaks of the two quantities are almost perfectly aligned, which we consider a promising counterexample to the statement that mean-field dynamics cannot address the above-quoted “fundamental problem” of quantum annealing. To substantiate this observation, we provide disorder-averaged statistics in the next section. To briefly comment on the inset of Fig. 7, which shows the outcomes of numerically exact simulations of the evolution under Eq.(2) for various final times $T_{f}$, we see that the ground state only begins to become visible around $T_{f}\sim 10^{4}$, in agreement with the expectation from the adiabatic theorem [6].
@@ III.2 Main Result: Disorder-Averaged Statistics
To go beyond the anecdotal evidence presented in the previous section, we now extend the analysis to an extended data set of small-minigap instances, the details of which are set out in Appendix B. The system and ensemble sizes of this data set are summarized in Tab. 2. These instances are obtained by computing the numerically exact spectrum and retaining only those that have a minigap $\Delta<\Delta_{\rm small}=10^{-2}$. As mentioned in the introduction, even among these instances there occur not a few that are solved exactly by adiabatic mean-field dynamics. We emphasize that we exclude these instances from the following analysis because the whole point is to devise an approximate method applicable to hard instances that cannot be solved in mean-field. The resulting effective ensemble sizes are given in Tab. 1. These remaining instances span a range of criticality as quantified by the peak of the corresponding adiabatic ratios.
For each instance in these ensembles, we determine the critical point $s_{*}$ according to Eq. (3) and then plot the corresponding maximally frustrated localization susceptibility against $s-s_{*}$, i.e. $\chi_{i_{*}}(s)$ must peak around the origin in order to have its maximum coincide with the adiabatic bottleneck. Importantly, we perform the disorder average of the random realizations of the SK spin glass after this transformation. The results are shown in Fig. 8: except for the smallest systems at $N=8$, the average localization susceptibility does indeed peak at the origin, which confirms that $\chi_{i_{*}}(s)$ indeed contains statistically relevant information about the onset of the adiabatic bottleneck. Note that if one only considers the local EA order parameter Eq. (11), the averaged agreement with the peak of the adiabatic ratio is worse than when taking the product with the maximally frustrated localization susceptibility. The same is true if considering only the average of the latter.
@@
Table 1: Effective ensemble size after excluding small-minigap instances solved by adiabatic mean-field evolution.
@@ IV Discussion & Conclusion
We have developed the theory of the nonequilibrium dynamics of the semi-classical spin coherent-state path integral, as encapsulated by Eqs. (9) and (II.4). This has allowed us to study the instance-specific onset of the adiabatic bottleneck for the quantum SK model via critical fluctuations. Our main result, Fig. 8, provides numerical evidence that this is statistically robust and not dependent on the system size.
It seems clear that there are many instances where the minigap with the first excited state is so small that one cannot hope of mitigating it practically in the near future. However, insofar as a bulk glass phase with a “soft” excitation spectrum [17] can be considered universal for long-range interacting spin glasses, there is the real possibility of doing even worse by scattering into higher excited states. This was indeed the case for our example instance above, and it seems to become more likely at larger $N$ (as indicated by Fig. 9 in Appendix B). A minimal intermediate goal could therefore be to mitigate these higher excitations, a goal for which the knowledge of the critical fluctuations, as presented in this work, appears to be very helpful.

### v2

@@ Abstract
The correspondence between long-range interacting quantum spin glasses and combinatorial optimization problems underpins the physical motivation for adiabatic quantum computing. On one hand, in disordered (quantum) spin systems, the focus is on exact methods such as the replica trick that allow the calculation of system quantities in the limit of infinite system and ensemble size. On the other hand, when solving a given instance of an optimization problem, disorder-averaged quantities are of no relevance, as one is solely interested in instance-specific, finite-size properties, in particular the optimal solution. Here, we apply the nonequilibrium Green-function formalism to the spin coherent-state path integral to obtain the statistical fluctuations and the collective-excitation spectrum along the annealing path. For the example of the quantum Sherrington-Kirkpatrick spin glass, by comparing to extensive numerically exact results, we show that this method provides access to the instance-specific bottlenecks of the annealing protocol.
@@ I Introduction
This leads us back to the “most fundamental problem” of quantum annealing [4]: how to obtain some information about the size and, in particular, the onset of the minigap without solving the problem. Here, we demonstrate that the semi-classical fluctuations around the spin mean-field introduced previously [10] can serve as an indicator of the critical point. Physically, these fluctuations describe the dynamics of paramagnons away from equilibrium. They are obtained by solving the equations of motion (22) of the nonequilibrium Green functions [11, 12], which in turn are derived from the Gaussian approximation of the spin coherent-state path integral [13, 11, 10].
@@
To obtain numerically exact data as a comparison for our semi-classical methods, we perform very extensive simulations. Specifically, we diagonalize a large number of random SK instances at different system sizes and retain only the subset with the smallest minimal spectral gaps. The instances in this subset then all contain bottlenecks of varying criticality when annealing the transverse field. The details of these data are provided in Appendix C. Interestingly, even within this critical subset, there remain instances for which the adiabatic mean-field evolution actually finds the true ground state. Our main analysis, however, is then further restricted to those critical instances for which the mean-field evolution fails. This appears natural since we want to investigate problems hard for quantum annealing – but any instance solved via mean-field can be considered easy. Importantly, we find that the information provided by our semi-classical approximation about the critical point of the annealing schedule is robust: the method works even when the mean-field algorithm fails, such that we succeed in showing that the alignment between the critical fluctuations and the adiabatic bottleneck is preserved also on average (Fig. 8).
@@ II Methods
The couplings $J_{ij}$ and the local magnetic fields $h_{i}$ are i.i.d. standard normal random variables (s. Appendix C). We set $\Gamma=1$ everywhere, i.e. we consider frequencies to be given in units of $\Gamma$. The adiabatic time is confined to $0\leq t\leq T_{f}$, which makes it natural to also introduce the scaled time $s(t)=t/T_{f}$, $s(t)\in[0,1]$. This results in the adiabatic Hamiltonian
@@ II.1 Mean-Field Dynamics
For very slow evolution, one finds that the trajectories following from Eqs. (9) have the property
@@ II.2 Maximally Frustrated Spins
As described vividly by Mézard, Parisi, and Virasoro in the introduction to Ref. [1], frustration is an important and ubiquitous phenomenon. Also in the present context, the tug of opposing forces can have dramatic consequences, leading to strongly frustrated spins that cannot “decide” whether to point to the north or south pole of their Bloch spheres. We illustrate this in Fig. 4, where one spin fluctuates between the hemispheres around the equator. For our random data set detailed in Appendix C, this can be observed to be quite characteristic of hard instances going through a critical point or the finite-size analog of a first-order phase transition, the onset of which finally forces the frustrated spin either north or south.
@@ II.3 Gaussian Fluctuations
The Gaussian fluctuations are bosonic and thus implicitly define creation and annihilation operators $\hat{\eta}_{i}^{{\dagger}}$, $\hat{\eta}_{i}^{\phantom{\dagger}}$ for quasi-particles one may imagine as paramagnons. In terms of these operators, we can introduce the so-called greater and lesser Green functions [23, 24]
$G_{ij}^{>}(t,t^{\prime})$ $=-\mathrm{i}\langle\hat{\eta}_{i}^{\phantom{\dagger}}(t)\hat{\eta}_{j}^{{\dagger}}(t^{\prime})\rangle,$ (18a)
$G_{ij}^{<}(t,t^{\prime})$ $=-\mathrm{i}\langle\hat{\eta}_{j}^{{\dagger}}(t^{\prime})\hat{\eta}_{i}^{\phantom{\dagger}}(t)\rangle.$ (18b)
The diagonals of these two matrices have an intuitive meaning at equal times, i.e. for $t^{\prime}\to t$ (s. also Fig. 5). By letting $\mathrm{i}G_{ii}^{>}(t,t)=\mathcal{N}_{i}+1$ and $\mathrm{i}G_{ii}^{<}(t,t)=\mathcal{N}_{i}$, we can interpret $\mathcal{N}_{i}$ as the number of paramagnons (for a bosonic light field, this would just be the number of photons). The only difference between the on-site greater and lesser Green functions at equal times is then given by the zero-point contribution to $G_{ii}^{>}(t,t)$. By inspecting Eq. (15), we see that the corresponding anomalous Green functions proportional to $\langle\hat{\eta}_{i}^{\phantom{\dagger}}(t)\hat{\eta}_{j}^{\phantom{\dagger}}(t^{\prime})\rangle$ will also appear in the description of the problem. For completeness, note that the time-dependent Bogoliubov Hamiltonian corresponding to Eq. (15) can be written as
$\begin{split}\hat{H}(t)=\frac{1}{2}\sum_{i,j=1}^{N}\Big{[}&A_{ij}(t)\hat{\eta}_{i}^{\dagger}\hat{\eta}_{j}^{\phantom{\dagger}}+A_{ji}(t)\hat{\eta}_{j}^{\dagger}\hat{\eta}_{i}^{\phantom{\dagger}}\\[-5.69054pt] +&B_{ij}(t)\hat{\eta}_{i}^{\dagger}\hat{\eta}_{j}^{\dagger}+\bar{B}_{ji}(t)\hat{\eta}_{i}^{\phantom{\dagger}}\hat{\eta}_{j}^{\phantom{\dagger}}\Big{]}.\end{split}$ (19)
@@ II.4 Schwinger-Keldysh Formalism
$\bm{G}^{>}(t,t^{\prime})$ $=-\mathrm{i}\left\langle\begin{pmatrix}\eta^{-}(t)\\ \bar{\eta}^{-}(t)\end{pmatrix}\begin{pmatrix}\bar{\eta}^{+}(t^{\prime})&\eta^{+}(t^{\prime})\end{pmatrix}\right\rangle,$ (21a)
$\bm{G}^{<}(t,t^{\prime})$ $=-\mathrm{i}\left\langle\begin{pmatrix}\eta^{+}(t)\\ \bar{\eta}^{+}(t)\end{pmatrix}\begin{pmatrix}\bar{\eta}^{-}(t^{\prime})&\eta^{-}(t^{\prime})\end{pmatrix}\right\rangle.$ (21b)
@@ II.5 Statistical and Spectral Functions
Instead of working with the Green functions defined by Eqs. (18) or (21), respectively, it is conventional to transform to the following equivalent linear combinations,
@@
We will refer to $\bm{F}(t,t^{\prime})$ as the statistical function, while $\bm{\rho}(t,t^{\prime})$ is the so-called spectral function. Similarly to our above discussion of the equal-time functions, it now follows that $\mathrm{i}F_{ii}(t,t)=2\mathcal{N}_{i}+1$ is directly related to the paramagnon number. The spectral function, in turn, has the property $\bm{\rho}(t,t)=-\mathrm{i}\mathds{1}$, the origin of which lies in the bosonic commutation relation. An illustration of the behavior of this function away from the equal-time diagonal is provided in Fig. 5. The equation of motion of the statistical function can be derived easily by combining Eqs. (22) to yield
@@
Furthermore, the solution of the full nonequilibrium dynamics, as developed here, requires two further equations of motion, one for the statistical and one for the spectral function. The former can be derived from Eq. (22a) by adding the respective equations for $\bm{G}^{>}$ and $\bm{G}^{<}$; it is not relevant here since we only require the equal-time dynamics $\bm{F}(T,0)=\bm{F}(t,t)$, which, in the Gaussian case, decouples from the dynamics away from the forward time diagonal $t=t^{\prime}$ (s. Fig. 5). The remaining equation of motion for the spectral function can also be derived straightforwardly from Eqs. (22) and reads
@@ III.1 Illustration: Details of a Hard Instance
For our current example, the total EA order parameter from Eq. (12), plotted in blue in the second panel, does not show a very distinct increase around the critical point, which is due to the fact that the already localized spins have the largest contributions. This underlines the usefulness of focusing on the critical spin. Finally, consider the effective magnetization shown in the last panel of Fig. 3. Intriguingly, we perceive that our critical spin is completely frustrated during the entire evolution, an observation that we have found repeated throughout the full data set described in Appendix C. For this reason, we refer to the critical spin as maximally frustrated.
@@
to which we will refer as the localization susceptibility. It provides a joint measure of how localized and susceptible to fluctuations the maximally frustrated spin is at any given time. The resulting curve for our current example is presented in Fig. 7 alongside the adiabatic ratio $g_{1}(s)$ as defined in Eq. (3). The peaks of the two quantities are almost perfectly aligned, which we consider a promising counterexample to the statement that mean-field dynamics cannot address the above-quoted “fundamental problem” of quantum annealing. To substantiate this observation, we provide disorder-averaged statistics in the next section.
To briefly comment on the inset of Fig. 7, which shows the outcomes of numerically exact simulations of the evolution under Eq. (2) for various final times $T_{f}$, we see that the ground state only begins to become visible around $T_{f}\sim 10^{4}$, in agreement with the expectation from the adiabatic theorem [6].
Regarding the role of the $q_{i}(s)$ factor in Eq. (29), it is helpful to also consider the additional results presented in Appendix B, where we apply our method to Maximum 2-Satisfiability (Max. 2-Sat.). While the most frustrated mean-field trajectory of our present example does localize to one of the Bloch-sphere poles at the end of the evolution, this is not always the case, as can be seen from Fig. 9. For this hard Max. 2-Sat. instance, while we find that $q_{i_{*}}(1)=0$, i.e. the most frustrated spin does not localize eventually, our definition of the localization susceptibility remains unaltered and leads to the result shown in Fig. 10.
@@ III.2 Main Result: Disorder-Averaged Statistics
To go beyond the anecdotal evidence presented in the previous section, we now extend the analysis to an extended data set of small-minigap instances, the details of which are set out in Appendix C. The system and ensemble sizes of this data set are summarized in Tab. 2. These instances are obtained by computing the numerically exact spectrum and retaining only those that have a minigap $\Delta<\Delta_{\rm small}=10^{-2}$. As mentioned in the introduction, even among these instances there occur not a few that are solved exactly by adiabatic mean-field dynamics. We emphasize that we exclude these instances from the following analysis because the whole point is to devise an approximate method applicable to hard instances that cannot be solved in mean-field. The resulting effective ensemble sizes are given in Tab. 1. These remaining instances span a range of criticality as quantified by the peak of the corresponding adiabatic ratios.
For each instance in these ensembles, we determine the critical point $s_{*}$ according to Eq. (3) and then plot the corresponding maximally frustrated localization susceptibility against $s-s_{*}$, i.e. $\chi_{i_{*}}(s)$ must peak around the origin in order to have its maximum coincide with the adiabatic bottleneck. Importantly, we perform the disorder average of the random realizations of the SK spin glass after this transformation. The results are shown in Fig. 8: except for the smallest systems at $N=8$, the average localization susceptibility does indeed peak at the origin, which confirms that $\chi_{i_{*}}(s)$ contains statistically relevant information about the onset of the adiabatic bottleneck. Note that if one only considers the local EA order parameter Eq. (11), the averaged agreement with the peak of the adiabatic ratio is worse than when taking the product with the maximally frustrated localization susceptibility. The same is true if considering only the average of the latter. We remark that while the alignment between localization susceptibility and adiabatic bottleneck is not always as ideal as in Fig. 7, there is almost always good agreement (as is necessary to obtain the significant average agreement observed in Fig. 8).
@@
Avg. Hamming distance 5.98 7.02 7.87 9.12 10.00 11.04
Table 1: Effective ensemble sizes and average Hamming distances to the optimal solution after excluding small-minigap instances solved by adiabatic mean-field evolution.
@@ IV Discussion & Conclusion
We have developed the theory of the nonequilibrium dynamics of the semi-classical spin coherent-state path integral, as encapsulated by Eqs. (9) and (22). This has allowed us to study the instance-specific onset of the adiabatic bottleneck for the quantum SK model via critical fluctuations. Our main result, Fig. 8, provides numerical evidence that this is statistically robust and not dependent on the system size.
It seems clear that there are many instances where the minigap with the first excited state is so small that one cannot hope of mitigating it practically in the near future. However, insofar as a bulk glass phase with a “soft” excitation spectrum [17] can be considered universal for long-range interacting spin glasses, there is the real possibility of doing even worse by scattering into higher excited states. This was indeed the case for our example instance above, and it seems to become more likely at larger $N$ (as indicated by Fig. 11 in Appendix C). A minimal intermediate goal could therefore be to mitigate these higher excitations, a goal for which the knowledge of the critical fluctuations, as presented in this work, appears to be very helpful.